OpenStudy (moazzam07):

Ted tossed a coin several times and recorded the number of heads and tails in the following table: Coin Experiment Side of Coin Number of Tosses Heads 18 Tails 22 What is the experimental probability of getting heads?

OpenStudy (anonymous):

\[experimental ~probability=\frac{ number~of~heads }{ number~of~tosses }\]

OpenStudy (moazzam07):

Ted tossed a coin several times and recorded the number of heads and tails in the following table: Coin Experiment Side of Coin Number of Tosses Heads 18 Tails 22 What is the experimental probability of getting heads? 0.18 0.22 0.35 0.45

OpenStudy (moazzam07):

some one help

OpenStudy (anonymous):

^^

OpenStudy (moazzam07):

@KryoWolf help

OpenStudy (anonymous):

how many tosses were there total? how many of the tosses were heads?

OpenStudy (anonymous):

look at the difference of tosses between heads and tails and you'll see that it's 4, then subtract two from tails and add it to heads and you would see that they would both be 20, so the total would be 40, 50%. 2 would be 5%, so subtract that 5% from the heads' side and you get 45%

OpenStudy (apricot):

Add the number of heads+tails. Then take the number of heads and divide it by that number. 18+22=40, 40/18 = .45. Your answer is 45%.

OpenStudy (moazzam07):

ah ok thanks @Apricot

OpenStudy (moazzam07):

@KryoWolf thanks

OpenStudy (apricot):

18/40= .45, sorry I mistyped haha.

OpenStudy (moazzam07):

@peachpi thanks

OpenStudy (anonymous):

no problem